OpenStudy (anonymous):

Find (g*f)(3) f(x)=3x-2 g(x)=x^2 I plugged three in... and I get 63... yet, that is not any of the answer choices. a. 1 c. 49 b. -8 d. 52

geerky42 (geerky42):

\[(g\cdot f)(x) = g(x)\cdot f(x)\]

geerky42 (geerky42):

What is \(g(3)\)? \(f(3)\)?

OpenStudy (anonymous):

g(3) = 9. f(3) = 7

geerky42 (geerky42):

Oh, is question \((g\circ f)(x)\)?

OpenStudy (anonymous):

Yes.

geerky42 (geerky42):

Oh that's completely different.

geerky42 (geerky42):

\[(g\circ f)(x) = g(~f(x)~)\]

OpenStudy (anonymous):

You're kidding me... This has never even been taught to me...

geerky42 (geerky42):

I'm not kidding lol

OpenStudy (anonymous):

okay, so, f(x) = 7. which leads to g(7) then that comes out to be 49.

OpenStudy (anonymous):

Thanks mate.! @geerky42